starting from the definition of Helmholtz free energy (A) derive relations for its variation with
(1) temperature at a constant volume and

(2) volume at a constant temperature.
1
Expert's answer
2019-09-16T03:53:27-0400

Helmholz free energy is given by:

"A=H-TS\\where\ \\U=internal\ energy\\T=temprature\\S=entropy\\A=helmholz\ free\ energy"


"Or,\\\ \ \ \ \ \ \ dA=dU-d(TS)\\\ \ \ \ \ \ \ dA=TdS-PdV-(TdS+SdT)"

"\ \ \ \ \ \ \ dA=-PdV-SdT\ \ \ \ \ \ \ .....\ \ \ \ (a)\\"


"(2)\ At\ constant\ temprature:SdT=0\\So,\ dA=-PdV\ \ \\\\(1)\ At\ constant\ volume:\ PdV=0\\So,\ dA=-SdT"